OpenStudy (anonymous):

Let h = ((v0^2)/4.9) sinθcosθ model the horizontal distance in meters traveled by a projectile. If the initial velocity is 44 meters/second, which equation would you use to find the angle needed to travel 150 meters?

OpenStudy (anonymous):

h = ((v(subscript o)^2)/4.9) sinθcosθ

OpenStudy (phi):

is this multiple choice or can do you just write down an equation ?

OpenStudy (anonymous):

A. 395.10sin(2θ) = 150 B. 150sin(2θ) = 150 C. 8.98sin(2θ) = 150 D. 197.55sin(2θ) = 150

OpenStudy (phi):

to do this problem you need to know \[ 2 \sin(\theta) \cos(\theta)= \sin(2\theta) \] if you divide both sides by 2, this is the same as \[\sin(\theta) \cos(\theta)= \frac{1}{2} \sin(2\theta) \]

OpenStudy (anonymous):

I was thinking C. 8.98sin(2θ) = 150

OpenStudy (phi):

the way to do it is start with \[ \frac{v_0^2}{4.9} \sinθ\cosθ = h\] now use the identity posted up above to replace sinθ cosθ also, replace v0 with 44 can you do that ?

OpenStudy (anonymous):

Oh.. Damn okay! Nevermind then. I got D now